1) We are in two good, two country, where all of the assumptions of the Heckscher-Ohlin model of trade are satisfied. Suppose that in Japan there are 50 workers and 50 acres of land, while in the US there are 100 workers and 200 acres of land. Suppose furthermore that the two goods, rice, and cloth, are produced using land and labor. Rice is land intensive. Suppose the two countries open up to trade. What will be the pattern of trade? why?
Answer #1

It has been provided that the Japan has 50 workers and 50 acres of land.

Worker-land ratio = Number of workers/Amount of land = 50/50 = 1

Land-worker ratio = Amount of land/Number of workers = 50/50 = 1

It has been provided that the US has 100 workers and 200 acres of land.

Worker-land ratio = Number of workers/Amount of land = 100/200 = 0.5

Land-worker ratio = Amount of land/Number of workers = 200/100 = 2

It can be seen that the Japan has greater Worker-land ratio than US. So, Japan is labor intensive country relative to US.

It can be seen that the US has greater Land-worker ratio than Japan. So, US is land intensive country relative to US.

According to the Heckscher-Ohlin model of trade, a country exports that good which it can produce using its intensive factor.

Since, rice is land intensive good and US is land intensive country, US will produce the rice and export it.

On the other hand, cloth is labor intensive and Japan is labor intensive country, Japan will produce cloth and export it.

So,

If the two countries open up to trade then,

Japan will export cloth and will import rice while US will export rice and will import cloth.